OpenStudy (anonymous):

Solve the differential equation y" + 4y' + 5 = 0

OpenStudy (anonymous):

let y=e^(mx) be the trial sol. then y'' = m^2* e^(mx) and y'= m*e^(mx) putting these in the eq. we have the auxillary sol. as m^2+4m +5= 0 (m+2)^2+1=0 hence m= -2 +i,-2-i hence the reqd sol. is y=e^(-2x)(Acosx+Bsinx) where A and B are constants